OpenStudy (anonymous):

Survey Question issue. How to make results go out of 100% total? 12 votes. Strawberry: 4 people liked. Chocolate: 3 people liked. Banana: 5 people liked. Thanks!

OpenStudy (anonymous):

I need to know what these would look like as percentages? This is for a project, please reply.

OpenStudy (anonymous):

o.o i wish i could hep but idk

OpenStudy (anonymous):

I'm not alone :D lol thanks anyways ^_^;; I wish one of those really smarty peoples would answer, but no one familiar is online...

terenzreignz (terenzreignz):

It looks like you're converting fractions to percent, am I right?

OpenStudy (anonymous):

Just results i got. Like, i asked what milkshake flavor people liked best...I have to covert these numbers to percentages that total 100% and im lost...

terenzreignz (terenzreignz):

Okay. Let's take chocolate, for instance. 3 out of your 12 respondents picked chocolate, yes?

OpenStudy (anonymous):

yup ^^

terenzreignz (terenzreignz):

Okay, so the fraction is \[\frac{3}{12}=\frac{1}{4}\]right?

OpenStudy (anonymous):

I think so :)

terenzreignz (terenzreignz):

More confidence ;) Anyway, to convert a fraction to percent, simply multiply it by 100 \[\huge \frac{1}{4}\times 100 = 25\] And then affix the percent sign so 3/12 = 25%

OpenStudy (anonymous):

ohhh!! :D i see now. Thank you!! :)

terenzreignz (terenzreignz):

No problem :)